# Positive polytopes with few facets in the Grassmannian

### v1

@@ 1 Introduction
Standard examples of positive geometries include polytopes in the projective space $\mathbb{P}^{n}$, the positive Grassmannian $\mathrm{Gr}_{\geq 0}(k,n)$ inside $\mathrm{Gr}(k,n)$, positive parts of toric varieties [1, Section 5], and the positive part $(\mathcal{M}_{0,n})_{\geq 0}$ of the Deligne-Knudsen-Mumford compactification $\overline{\mathcal{M}_{0,n}}$ of the moduli space of rational curves with $n$ marked points [2, Proposition 8.2]. All these objects have one thing in common: a remarkably rich underlying combinatorial structure. In particular, all of them allow for a stratification of their algebraic boundary (e.g. the face stratification for polytopes and the positroid stratification for $\mathrm{Gr}_{\geq 0}(k,n)$ [15]). The fact that their boundaries admit such a stratification allows to approach the question of whether these objects are positive geometries (and finding their canonical forms) from quite a combinatorial point of view, via studying their residual arrangements. For a set whose boundary has a well defined stratification consisting of faces contained in the intersection of a finite set of facets, there is a residual arrangement. It consists of all irreducible components of intersections of (Zariski closures of) facets that are not Zariski closures of a face. The canonical form has simple poles along each facet, so the denominator of the rational function defining the canonical form is the product of the defining equations of the facets. The numerator defines a hypersurface of minimal possible degree interpolating the residual arrangement. Hypersurfaces with this property are called adjoint. This classical algebraic geometric term denotes a divisor whose strict transform on the blow up along the residual arrangement restricts to a canonical divisor on the strict transform of the algebraic boundary. The uniqueness of the canonical form may be achieved by the uniqueness of such a canonical divisor, i.e. by the uniqueness of the adjoint.
@@ 2 Preliminaries
Let $P\subseteq\mathbb{P}^{n}$ be a polytope. The collection of intersections in $\mathbb{P}^{n}$ of hyperplanes in $\mathcal{H}_{P}$ that do not contain any face of $P$ of codimension at least two is called the residual arrangement of $P$ and is denoted by $\mathcal{R}(P)$.
@@
Let $P\subseteq\mathbb{P}^{n}$ be a polytope with $d$ facet hyperplanes. A hypersurface in $\mathbb{P}^{n}$ of degree $d-n-1$ containing $\mathcal{R}(P)$ is called an adjoint hypersurface of $P$. The defining polynomial of an adjoint hypersurface is called an adjoint polynomial.
@@
Adjoint hypersurfaces play an important role in the context of positive geometries [1], where they arise as numerators of canonical forms of polytopes. We now define positive geometries and canonical forms, following [12]. Let $X$ be a complex $d$-dimensional irreducible algebraic variety defined over $\mathbb{R}$. We equip the real points $X(\mathbb{R})$ with the analytic topology. Let $X_{\geq 0}\subset X(\mathbb{R})$ be a closed semialgebraic subset such that the interior $X_{>0}=\operatorname{Int}(X_{\geq 0})$ is an oriented $d$-manifold, and the closure of $X_{>0}$ recovers $X_{\geq 0}$. Let $\partial X_{\geq 0}$ denote the boundary $X_{\geq 0}\setminus X_{>0}$ and let $\partial X$ denote the algebraic boundary of $X$, i.e. the Zariski closure of $\partial X_{\geq 0}$. Let $C_{1},C_{2},\ldots,C_{r}$ be the irreducible components of $\partial X$. Let $C_{i,\geq 0}$ denote the closures of the interior of $C_{i}\cap X_{\geq 0}$ in $C_{i}(\mathbb{R})$. The spaces $C_{1,\geq 0},\ldots,C_{r,\geq 0}$ are called the boundary components, or facets of $X_{\geq 0}$.
@@
Polytopes in $\mathbb{P}^{n}$ are known to be positive geometries [1, Section 5], and the canonical form of a polytope is given by $f(\mathbf{x})\mu_{\mathbb{P}^{n}}(\mathbf{x})$, where $f(\mathbf{x})$ is a rational function, $\mu_{\mathbb{P}^{n}}(\mathbf{x})=\sum_{i=0}^{n}x_{i}dx_{0}\wedge\ldots\wedge\widehat{dx_{i}}\wedge\ldots\wedge dx_{n}$ is the standard rational top-form on $\mathbb{P}^{n}$ and $\mathbf{x}=(x_{0}:\ldots:x_{n})$ is the vector of homogeneous coordinates. The numerator of $f$ is then given by the adjoint polynomial, and the denominator is the product of linear forms defining the facets.
@@
The real Grassmannian $\mathrm{Gr}(k,n)$ is the algebraic variety parametrizing $k$-dimensional linear subspaces of $\mathbb{R}^{n}$ (respectively $(k-1)$-dimensional subspaces of $\mathbb{P}^{n-1}$). The Grassmannian $\mathrm{Gr}(k,n)$ can be embedded into the projective space $\mathbb{P}^{\binom{n}{k}-1}$ via the Plücker embedding. This is done by sending a $k\times n$ matrix whose rows span a vector space to the vector of its maximal minors, called the Plücker coordinates. See e.g. [13, Chapter 5] for an introductory reference.
@@
By setting some of the Plücker coordinates to zero, we get parts of the boundary of $\mathrm{Gr}_{\geq 0}(k,n)$ called positroid cells. The unique top-dimensional cell, defined by requiring all Plücker coordinates to be nonzero, is the interior of $\mathrm{Gr}_{\geq 0}(k,n)$. Zariski closures of positroid cells are called positroid varieties.
It is the stratification into positroid cells that defines the structure of a regular CW-complex on $\mathrm{Gr}_{\geq 0}(k,n)$. In this context, positroid cells are analogs of faces of a polytope. This allows to define the residual arrangement of $\mathrm{Gr}_{\geq 0}(k,n)$ as the collection of irreducible components of intersections of positroid varieties of codimension one that are not contained in $\mathrm{Gr}_{\geq 0}(k,n)$ itself. This definition is unfortunately not particularly meaningful: it is known that $\mathrm{Gr}_{\geq 0}(k,n)$ is a simplex-like positive geometry [1, Section 5.5.2], that is, its residual arrangement is empty (in analogy with that of the simplex in $\mathbb{P}^{n}$). This similarity between the positive Grassmannian and simplices stems from the fact that $\mathrm{Gr}_{\geq 0}(k,n)$ is the intersection of the whole Grassmannian $\mathrm{Gr}(k,n)$ with the positive simplex $\mathbb{P}^{\binom{n}{k}-1}_{\geq 0}$. However, when $\mathrm{Gr}(k,n)$ is intersected with more complicated polytopes in $\mathbb{P}^{\binom{n}{k}-1}_{\geq 0}$, the resulting object may have non-empty residual arrangement. This paper studies a special case of this situation. A natural question to ask is whether there is a unique adjoint hypersurface, and if not, what is the dimension of the family of adjoint polynomials.
@@ 3 Positive polytopes in $\mathrm{Gr}(2,4)$
We adopt the notation $\mathbf{p}=(p_{12},p_{13},p_{14},p_{23},p_{24},p_{34})$ The positive simplex $\mathbb{P}^{5}_{\geq 0}$ in $\mathbb{P}^{5}$ is defined by the inequalities $p_{ij}\geq 0$ for $1\leq i<j\leq 4$. Note that each of these equalities does not make sense individually, since the Plücker coordinates are only defined up to a common scalar. What we mean by these six linear inequalities is that the quadratic inequalities $p_{ij}p_{kl}\geq 0$ hold for any pair of Plücker coordinates. In a similar way, in what follows, whenever we define a set by linear inequalities $l_{i}(\mathbf{p})\geq 0$ in the projective space, we will mean the set defined by the corresponding quadratic inequalities obtained by multiplying pairs of linear forms $l_{i}(\mathbf{p})$ in all possible ways.
@@
Let $P\subset\mathbb{P}^{5}$ be a simple polytope with $m+6$ facets such that six of the facets are contained in the Plücker hyperplanes and the remaining facets are in the hyperplanes $h_{i}(\mathbf{p})=0$, $i=1,\ldots,m$. Assuming that the semialgebraic set $S=P\cap\mathrm{Gr}(2,4)$ has $m+4$ facets that are given by the four positroid hyperplanes $p_{12}=0$, $p_{14}=0$, $p_{23}=0$, $p_{34}=0$ and the $m$ additional hyperplanes $h_{i}(\mathbf{p})=0$, we call $S$ a positive polytope in $\mathrm{Gr}(2,4)$ and $P$ the associated polytope of $S$.
@@
Let $S$ be a positive polytope in $\mathrm{Gr}(2,4)$. Just like in the case of polytopes in $\mathbb{P}^{5}$, the residual arrangement of $S$ consists of intersections of Zariski closures of facets of $S$ that do not contain a face of $S$.
@@ 3.1 Combinatorial classification of positive hexahedra
The associated polytope of a positive hexahedron in $\mathrm{Gr}(2,4)$ is a simple 5-dimensional polytope with $8$ facets. There are 8 combinatorial types of such polytopes (they are labeled by Gale diagrams of their polar duals in [9, Figure 6.3.3]). The combinatorial type of the associated polytope determines the number of linear spaces of each dimension in its residual arrangement, as well as the incidence relations between these spaces. However, when intersected with $\mathrm{Gr}(2,4)$, two combinatorially equivalent associated polytopes can produce positive hexahedra with meaningfully different residual arrangements (see Example 3.4). For this reason, we need a finer notion of combinatorial type for positive hexahedra that would capture the presence of $\mathrm{Gr}(2,4)$. We now introduce this.
@@
Given a polytope in $\mathbb{P}^{5}$ with $8$ facets, one can choose coordinates on $\mathbb{P}^{5}$ such that any given ordered subset of six facets is given by the vanishing of the Plücker coordinates. Since the Grassmannian is given by the quadratic equation $p_{12}p_{34}-p_{13}p_{24}+p_{14}p_{23}=0$, to determine the residual arrangement of a positive hexahedron $S$, one needs to know the combinatorial type of the associated polytope $P$ and which intersections of two facet hyperplanes are given by the vanishing of the Plücker monomials $p_{12}p_{34}$, $p_{13}p_{24}$ and $p_{14}p_{23}$. We say that two positive hexahedra $S_{1}$ and $S_{2}$ with associated polytopes $P_{1}$ and $P_{2}$ defined by additional hyperplanes $h_{1}(\mathbf{p}),h_{2}(\mathbf{p})$ and $h_{1}^{\prime}(\mathbf{p}),h_{2}^{\prime}(\mathbf{p})$ respectively are combinatorially equivalent if $P_{2}$ can be obtained from $P_{1}$ by replacing $h_{1}$ and $h_{2}$ with $h_{1}^{\prime}$ and $h_{2}^{\prime}$ (the order is not important) and permuting the Plücker coordinates in a way that preserves the set of Plücker monomials.
@@
By definition above, a choice of combinatorial type of $S$ is a choice of combinatorial type of $P$ along with the choice of three intersections of hyperplanes defined by the Plücker monomials. There are $420$ ways to choose three intersections of two out of eight facet hyperplanes. For each combinatorial type of the associated polytope this yields $420$ combinatorial types of positive hexahedra in $\mathrm{Gr}(2,4)$. The total number of combinatorial types of positive hexaherda is therefore $8\cdot 420=3360$.

### v2

@@ 1 Introduction
Standard examples of positive geometries include polytopes in the projective space $\mathbb{P}^{n}$, the positive Grassmannian $\mathrm{Gr}_{\geq 0}(k,n)$ inside $\mathrm{Gr}(k,n)$, positive parts of toric varieties [1, Section 5], and the positive part $(\mathcal{M}_{0,n})_{\geq 0}$ of the Deligne-Knudsen-Mumford compactification $\overline{\mathcal{M}_{0,n}}$ of the moduli space of rational curves with $n$ marked points [2, Proposition 8.2]. All these objects have one thing in common: a remarkably rich underlying combinatorial structure. In particular, all of them allow for a stratification of their algebraic boundary (e.g. the face stratification for polytopes and the positroid stratification for $\mathrm{Gr}_{\geq 0}(k,n)$ [15]). The fact that their boundaries admit such a stratification allows to determine whether these objects are positive geometries and find their canonical forms by using a combinatorial approach, namely by studying their residual arrangements. While we refer the reader to Section 2 for precise definitions, we will now briefly explain the idea of this approach. For a set whose boundary has a well defined stratification consisting of faces contained in the intersection of a finite set of facets, there is a residual arrangement. It consists of all irreducible components of intersections of (Zariski closures of) facets that are not Zariski closures of a face. The canonical form has simple poles along each facet, so the denominator of the rational function defining the canonical form is the product of the defining equations of the facets. The canonical form is only allowed to have poles along the faces, and therefore the numerator of the rational function must cancel the zeros of the denominator located away from the faces of our set. These zeros are precisely in the residual arrangement. Thus, the numerator defines a hypersurface of minimal possible degree interpolating the residual arrangement. Hypersurfaces with this property are called adjoint. This classical algebraic geometric term denotes a divisor whose strict transform on the blow up along the residual arrangement restricts to a canonical divisor on the strict transform of the algebraic boundary. The uniqueness of the canonical form may be achieved by the uniqueness of such a canonical divisor, i.e. by the uniqueness of the adjoint.
@@ 2 Preliminaries
Let $P\subseteq\mathbb{P}^{n}$ be a polytope. The collection of all intersections of hyperplanes in $\mathcal{H}_{P}$ that do not contain a face of $P$ is called the residual arrangement of $P$ and is denoted by $\mathcal{R}(P)$.
@@
Let $P\subseteq\mathbb{P}^{n}$ be a polytope with $d$ facet hyperplanes. A hypersurface in $\mathbb{P}^{n}$ of degree dn1d-n-1 containing $\mathcal{R}(P)$ is called an adjoint hypersurface of $P$. The defining polynomial of an adjoint hypersurface is called an adjoint polynomial.
@@
Adjoint hypersurfaces play an important role in the context of positive geometries [1], where their defining polynomials arise as numerators of canonical forms of polytopes. We now define positive geometries and canonical forms, following [12]. Let $X$ be a complex $d$-dimensional irreducible algebraic variety defined over $\mathbb{R}$. We equip the real points $X(\mathbb{R})$ with the analytic topology. Let $X_{\geq 0}\subset X(\mathbb{R})$ be a closed semialgebraic subset such that the interior $X_{>0}=\operatorname{Int}(X_{\geq 0})$ is an oriented $d$-manifold, and the closure of $X_{>0}$ recovers $X_{\geq 0}$. Let $\partial X_{\geq 0}$ denote the boundary $X_{\geq 0}\setminus X_{>0}$ and let $\partial X$ denote the algebraic boundary of $X$, i.e. the Zariski closure of $\partial X_{\geq 0}$. Let $C_{1},C_{2},\ldots,C_{r}$ be the irreducible components of $\partial X$. Let $C_{i,\geq 0}$ denote the closures of the interior of $C_{i}\cap X_{\geq 0}$ in $C_{i}(\mathbb{R})$. The spaces $C_{1,\geq 0},\ldots,C_{r,\geq 0}$ are called the boundary components, or facets of $X_{\geq 0}$. In Figure 1 the algebraic boundary of the pentagon consists of the five black lines. The components $C_{i}$ are the individual lines, and the sets $C_{i,\geq 0}$ are the edges of the pentagon.
@@
Polytopes in $\mathbb{P}^{n}$ are known to be positive geometries [1, Section 5], and the canonical form of a polytope is given by $f(\mathbf{x})\mu_{\mathbb{P}^{n}}(\mathbf{x})$, where $f(\mathbf{x})$ is a rational function, $\mu_{\mathbb{P}^{n}}(\mathbf{x})=\sum_{i=0}^{n}x_{i}dx_{0}\wedge\ldots\wedge\widehat{dx_{i}}\wedge\ldots\wedge dx_{n}$ is the standard rational top-form on $\mathbb{P}^{n}$ and $\mathbf{x}=(x_{0}:\ldots:x_{n})$ is the vector of homogeneous coordinates. The numerator of $f$ is then given by the adjoint polynomial, and the denominator is the product of linear forms defining the facets [12, Theorem 5].
@@
The real Grassmannian $\mathrm{Gr}(k,n)$ is the algebraic variety parametrizing $k$-dimensional linear subspaces of $\mathbb{R}^{n}$ (or equivalently, (k1)(k-1)-dimensional subspaces of $\mathbb{P}^{n-1}$). The Grassmannian $\mathrm{Gr}(k,n)$ can be embedded into the projective space $\mathbb{P}^{\binom{n}{k}-1}$ via the Plücker embedding. This is done by sending a $k\times n$ matrix whose rows span a vector space to the vector of its maximal minors, called the Plücker coordinates. See e.g. [13, Chapter 5] for an introductory reference.
@@
The boundary strata of $\mathrm{Gr}_{\geq 0}(k,n)$ are called (non-top-dimensional) positroid cells. They are defined by intersecting $\mathrm{Gr}_{\geq 0}(k,n)$ with linear spaces given by the vanishing of some of the Plücker coordinates. The unique top-dimensional positroid cell, defined by requiring all Plücker coordinates to be nonzero, is the interior of $\mathrm{Gr}_{\geq 0}(k,n)$. Zariski closures of positroid cells are called positroid varieties.
It is the stratification into positroid cells that defines the structure of a regular CW-complex on $\mathrm{Gr}_{\geq 0}(k,n)$. In this context, positroid cells are analogs of faces of a polytope. This allows one to define the residual arrangement of $\mathrm{Gr}_{\geq 0}(k,n)$ as the collection of irreducible components of intersections of positroid hypersurfaces in $\mathrm{Gr}(k,n)$ that are not contained in $\mathrm{Gr}_{\geq 0}(k,n)$ itself. This definition is unfortunately not particularly meaningful: it is known that $\mathrm{Gr}_{\geq 0}(k,n)$ is a simplex-like positive geometry [1, Section 5.5.2], that is, its residual arrangement is empty (in analogy with that of the simplex in $\mathbb{P}^{n}$). This similarity between the positive Grassmannian and simplices stems from the fact that $\mathrm{Gr}_{\geq 0}(k,n)$ is the intersection of the whole Grassmannian $\mathrm{Gr}(k,n)$ with the positive simplex $\mathbb{P}^{\binom{n}{k}-1}_{\geq 0}$. However, when $\mathrm{Gr}(k,n)$ is intersected with more complicated polytopes in $\mathbb{P}^{\binom{n}{k}-1}_{\geq 0}$, the resulting object may have non-empty residual arrangement. This paper studies a special case of this situation. Natural questions to ask are whether there is a unique adjoint hypersurface, and if not, what the dimension of the family of adjoint polynomials is.
@@ 3 Positive polytopes in $\mathrm{Gr}(2,4)$
We adopt the notation $\mathbf{p}=(p_{12},p_{13},p_{23},p_{14},p_{24},p_{34})$ . The positive simplex $\mathbb{P}^{5}_{\geq 0}$ in $\mathbb{P}^{5}$ is defined by the inequalities $p_{ij}\geq 0$ for $1\leq i<j\leq 4$. Note that each of these equalities does not make sense individually, since the Plücker coordinates are only defined up to a common scalar. What we mean by these six linear inequalities is that the quadratic inequalities $p_{ij}p_{kl}\geq 0$ hold for any pair of Plücker coordinates. In a similar way, in what follows, whenever we define a set by linear inequalities $l_{i}(\mathbf{p})\geq 0$ in the projective space, we will mean the set defined by the corresponding quadratic inequalities obtained by multiplying pairs of linear forms $l_{i}(\mathbf{p})$ in all possible ways.
@@
Let $P\subset\mathbb{P}^{5}$ be a simple polytope with $m+6$ facets such that six of the facets are contained in the Plücker hyperplanes and the remaining facets are in the hyperplanes defined by the vanishing of the linear forms $h_{i}(\mathbf{p})=0$, $i=1,\ldots,m$. Assuming that the semialgebraic set $S=P\cap\mathrm{Gr}(2,4)$ has $m+4$ facets that are given by the four positroid hyperplanes $p_{12}=0$, $p_{14}=0$, $p_{23}=0$, $p_{34}=0$ and the $m$ additional hyperplanes $h_{i}(\mathbf{p})=0$, we call $S$ a positive polytope in $\mathrm{Gr}(2,4)$ and $P$ the associated polytope of $S$.
@@
Let $S$ be a positive polytope in $\mathrm{Gr}(2,4)$. Just like in the case of polytopes in $\mathbb{P}^{5}$, the residual arrangement of $S$ consists of those irreducible components of intersections of Zariski closures of facets of $S$ that do not contain a face of $S$.
@@ 3.1 Combinatorial classification of positive hexahedra
The associated polytope of a positive hexahedron in $\mathrm{Gr}(2,4)$ is a simple five-dimensional polytope with eight facets. There are eight combinatorial types of such polytopes (they are labeled by Gale diagrams of their polar duals in [9, Figure 6.3.3]). The combinatorial type of the associated polytope determines the number of linear spaces of each dimension in its residual arrangement, as well as the incidence relations between these spaces. However, when intersected with $\mathrm{Gr}(2,4)$, two combinatorially equivalent associated polytopes can produce positive hexahedra with meaningfully different residual arrangements (see Example 3.4). For this reason, we need a finer notion of combinatorial type for positive hexahedra that would capture the presence of $\mathrm{Gr}(2,4)$. We now introduce this.
@@
Given a polytope in $\mathbb{P}^{5}$ with eight facets, one can choose coordinates on $\mathbb{P}^{5}$ such that any given ordered subset of six facets is given by the vanishing of the Plücker coordinates. Since the Grassmannian is given by the quadratic equation $p_{12}p_{34}-p_{13}p_{24}+p_{14}p_{23}=0$, to determine the residual arrangement of a positive hexahedron $S$, one needs to know the combinatorial type of the associated polytope $P$ and which intersections of two facet hyperplanes are given by the vanishings of the Plücker coordinates $p_{12}=p_{34}=0$, $p_{13}=p_{24}=0$ and $p_{14}=p_{23}=0$. We now elaborate on this. Suppose an octahedron $P$ in $\mathbb{P}^{5}$ is defined by the facet inequalities $h_{i}(\mathbf{p})\geq 0,\ i=1,\ldots 8$. Let $I\subseteq[8]$. To determine whether the intersection of facet hyperplanes $\{h_{i}(\mathbf{p})=0\ :\ i\in I\}$ is in the residual arrangement $\mathcal{R}(P)$, one needs to determine whether the system of equations and inequalities $h_{i}(\mathbf{p})=0,\ i\in I$ and $h_{j}(\mathbf{p})\geq 0,\ j\not\in I$ has a solution. Intersecting the polytope $P$ with a quadratic hypersurface $q(\mathbf{p}):=h_{i_{1}}(\mathbf{p})h_{i_{2}}(\mathbf{p})-h_{i_{3}}(\mathbf{p})h_{i_{4}}(\mathbf{p})+h_{i_{5}}(\mathbf{p})h_{i_{6}}(\mathbf{p})=0$ yields a semialgebraic set $S$. To determine whether the intersection of facets of this set $\{q(\mathbf{p})=0,\ h_{i}(\mathbf{p})=0\ :\ i\in I\}$ is in the residual arrangement $\mathcal{R}(S)$, one now needs to know which of the $h_{i}$’s participate in the equation for $q$. Note that permuting $i_{1}$ and $i_{2}$ (and analogously $i_{3}$ and $i_{4}$ as well as $i_{5}$ and $i_{6}$) does not change the equation of $q$. Moreover, simultaneously exchanging $i_{1}$ and $i_{5}$ and $i_{2}$ and $i_{6}$ does not change this equation. In our setup, the quadric $q$ is the Plücker quadric, and we need to choose which of the $h_{i}$’s define its three monomials.
We say that two positive hexahedra $S_{1}$ and $S_{2}$ with associated polytopes $P_{1}$ and $P_{2}$ defined by additional hyperplanes $h_{1}(\mathbf{p}),h_{2}(\mathbf{p})$ and $h_{1}^{\prime}(\mathbf{p}),h_{2}^{\prime}(\mathbf{p})$ respectively are combinatorially equivalent if $P_{2}$ can be obtained from $P_{1}$ by replacing $h_{1}$ and $h_{2}$ with $h_{1}^{\prime}$ and $h_{2}^{\prime}$ (the order is not important) and permuting the Plücker coordinates in a way that preserves the Plücker quadric $p_{12}p_{34}-p_{13}p_{24}+p_{14}p_{23}$.
@@
By the definition above, a choice of combinatorial type of $S$ is a choice of combinatorial type of $P$ along with the choice of three intersections of hyperplanes defined by the equations $p_{12}=p_{34}=0$, $p_{13}=p_{24}=0$ and $p_{14}=p_{23}=0$, in which we identify the choices that exchange the spaces $p_{12}=p_{34}=0$ and $p_{14}=p_{23}=0$. There are $420$ ways to choose three intersections of two out of eight facet hyperplanes, which come in pairs whose elements get identified as above. For each combinatorial type of the associated polytope this yields $210$ combinatorial types of positive hexahedra in $\mathrm{Gr}(2,4)$. The total number of combinatorial types of positive hexaherda is therefore $8\cdot 210=1680$.